# The importance of visual modelling languages in generative software engineering

### v2

@@ 2.1 Software Engineering and UML
The Unified Modelling Language (UML) emerged in 1990s as a response to the disparate notational systems and approaches to software design Rumbaugh et al. (2004) utilised across the phases of the software development process. UML is a general-purpose standardised visual language for designing systems. In SE, UML is utilised in three main areas: use case development, static analysis, and dynamic analysis of the software system. Use case development is a key step of requirement analysis. Use cases [^1] User stories in Agile Software Development Beck et al. (2001) describe how a user interacts with a system or product to achieve a specific goal. Static analysis of a software system describes the structure of a system by showing its classes, their attributes, operations (or methods), and the relationships among objects. Dynamic analysis expresses and model the behaviour of the system over time. More specifically, in UML, use case diagrams support use case development. Class diagrams support static analysis; and interaction (sequence, activity, collaboration) diagrams support dynamic analysis. However, it should be noted that UML is a vast language that finds countless applications to go beyond the three areas here considered.
@@ 2.2 Generative Artificial Intelligence
Prompts are instructions or input given to a large language model (LLM) to generate a specific response. Prompt engineering, that is the design prompting strategies to query language models, offers a cost-effective way to adapt pre-trained models without full fine-tuning. Prompt engineering techniques include: Single Prompt Techniques (e.g. Zero-Shot, Few-Shot), Multiple Prompt Techniques (e.g. Prompt Chaining), and the use of LLM frameworks with external tools (e.g. Retrieval-Augmented Generation Fan et al. (2024), and ReAct Yao et al. (2023)).
@@ 2.3 The role of GenAI in Software Engineering
Code generation has been object of investigation for a long time in the AI community. Early works used symbolic and neural-semiotic approaches Alur et al. (2013). However, recent neurolinguistic models, such as GPT-4 Liu et al. (2024) and Copilot Ma et al. (2023), can generate code directly from natural language descriptions. This capability has significantly advanced code completion, automatic code generation, and the transformation of natural language annotations into code, empowering developers with powerful tools and driving automation in software development. None of the studies listed in the above surveys focus on code generation leveraging multimodal prompts that include sketches & diagrams.
@@ 3 Methodology
Our approach takes the form of an exploratory case study Priya (2020), which aims to identify new research questions that can be investigated in future studies. More specifically, we develop a portfolio of novel GenSE use cases that, to the best of our knowledge, have not been previously investigated in the literature.
In our study, we utilise Microsoft Copilot in its web-based version, which is based on GPT-4o and allows image attachments. To ensure reproducibility of the discussion in Sections 4.1-4.5, we have also developed an equivalent pipeline in Gemini, formalised in a Jupyter Notebook based on gemini-1.5-flash, which is included in the supplementary material (SM). In both cases, we left all LLM parameters to their default settings and we did not specify any role or context instructions.
Table 1 maps use cases discussed in the rest of this work to relevant SE sub-tasks of interest. All use cases are based on an interaction with the LLMs that takes the form of an interactive chat comprising multiple rounds of questions and responses, which allow users to step incrementally toward answers and thus get help with multipart problems. The core principle underpinning all our use cases is to leverage UML diagrams in order to guide the software development process. More specifically, in Section 4.1 we leverage class diagrams to guide the LLM in the context of implementing relevant classes, attributes, and operations; in Section 4.2, we contrast diagram-based and natural language-based descriptions of a system; in Section 4.3 we leverage interaction diagrams to guide the LLM in the context of implementing the desired system behaviour; in Section 4.4 we leverage hand drawn activity diagrams to guide the LLM in the context of implementing a given method; in Section 4.5 we leverage design patterns to influence the design of a given system. Finally, in our SM we present three additional case studies: in the first case study, we leverage hand drawn diagrams and design patterns to implement a mathematical expressions evaluator; in the second and third case studies, which feature a higher degree of complexity, we leverage class diagrams to implement a tic tac toe game and a game of checkers, respectively.
@@
information transfer 4.2 SD
dynamic modelling 4.3 SD/DE/CS
hand drawn diagrams 4.4 DE/CS
design patterns 4.5 SD
expression evaluator SM SD/DE
tic tac toe SM SD/DE
checkers SM SD/DE
@@ 4 Multimodal GPTs in software development
The role of sketches and diagrams in the daily work of software developers has been investigated in Baltes and Diehl (2014). This study found that most practitioners produced informal hand-drawn diagrams and did not use UML; those using UML, tended to use it informally and selectively. We argue that this stems from the fact that the adoption of a “formal” notation, at present, bears no advantage with respect to an “informal” one. Over several decades, firms have repeatedly tried to develop tools (e.g. IBM Rational Rose) that could automatically generate working code from UML diagrams, or reverse engineer UML diagram from existing code. While these technologies still exist, it is rare to find developers who routinely develop a complete set of UML diagrams and then translate this to code using such tools. We find it plausible that the lacklustre success of these tools is due to a fundamental misunderstanding of what UML is: a language for capturing and communicating conceptual requirements, not one aimed at describing a complete system. In other words, UML diagrams are most useful when they are high level and sufficiently abstract. Generation of complete working code requires such diagrams to reach a level of detail that is equivalent to the working code itself; but generating diagrams at this level of abstraction would be a complete waste of time: why then not generating the code itself directly?
@@ 4.1 Static modelling
4.2 Comparison to natural language
We consider once more the class diagram in Figure 1. To illustrate the advantage of using diagrams over natural language, consider the following prompt.
Describe the attached class diagram via natural language.
The result returned by Copilot is illustrated in Figure 3, this description is 1375 characters long.
The attached class diagram represents a library system with the following classes and relationships:
  1. 1.
    Classes:
    • Book: Represents a book in the library.
    • Copy: Represents a copy of a book or journal.
    • Journal: Represents a journal in the library.
    • LibraryMember: Represents a member of the library.
    • MemberOfStaff: Represents a staff member of the library.
  2. 2.
    Relationships:
    • Book to Copy: A book has a one-to-many relationship with copies, indicated by ”1” on the Book side and ”1..*” on the Copy side. This means one book can have multiple copies.
    • LibraryMember to Copy: A library member can borrow or return copies, indicated by ”0..1” on the LibraryMember side and ”0..*” on the Copy side. This means a library member can borrow or return zero or one copy, and a copy can be borrowed or returned by zero or more library members.
    • MemberOfStaff to LibraryMember: A member of staff is a type of library member, indicated by the inheritance arrow pointing from MemberOfStaff to LibraryMember.
    • MemberOfStaff to Journal: A member of staff can borrow or return journals, indicated by ”0..1” on the MemberOfStaff side and ”0..*” on the Journal side. This means a member of staff can borrow or return zero or one journal, and a journal can be borrowed or returned by zero or more staff members.
Figure 3: The textual description of the class diagram returned by Copilot
Contrast this with the PlantUML notation description of the class diagram illustrated in Figure 4, which is 302 characters long, that is only 20% of the size of the natural language description.
@startuml
class Book {
class Copy {
class Journal {
class LibraryMember {
class MemberOfStaff {
LibraryMember <|-- MemberOfStaff
LibraryMember "0..1" -- "0..*" Copy : borrows/returns
LibraryMember "0..1" -- "0..*" Journal : borrows/returns
Copy "1" -- "1..*" Book : is a copy of
@enduml
Figure 4: The description of the class diagram in PlantUML notation
This example makes apparent the point discussed in the paragraph from Tufte (2003) we quoted in Section 1: the visual channel is an intensely high-resolution channel. Moreover, it reinforces the findings in Li et al. (2024) on the interrelation between understanding, learning, and compression. Further examples contrasting diagrams and natural language descriptions are provided in the SM.
@@ 4.3 Dynamic modelling
We may now want to proceed and describe the dynamic behaviour of the system. Rather than using natural language, we can describe specific aspect of such behaviour by leveraging suitable UML interaction diagrams, such as the sequence diagram in Figure 5.
@@
Unfortunately, the code obtained is not fully functional. There is a bug in method borrow_copy of class LibraryMember. This bug can easily be resolved by adding a single instruction (more details are provided in the SM). In line with contemporary SE and Agile practices — which predicate an iterative form of software development aimed at progressively discovering how to meet customer requirements — further iterations of the process, potentially involving additional diagrams and further prompts, can be used to obtain working code.
@@ 4.4 Hand-drawn diagrams
In this specific instance, Copilot returned a diagram with a small number of syntax errors. The errors were minor issues with the PlantUML syntax of the two while loops in the diagram, and could be easily fixed by hand. The resulting activity diagram is shown in Figure 11. Conversely, a second prompt asking Copilot to generate a diagram in PlantUML notation by translating directly the diagram in Figure 10 produced no errors. Once more, the focus of our discussion is facilitating, and not fully automating, SE activities. Therefore, the presence of minor errors is of no concern.
@@ 5 Discussion and conclusions
We shall focus first on the research gaps we bridged in GenSE. First, our study addresses the lack of applications of GenAI to the software design sub-task of SE, a literature gap identified in Huang et al. (2024); Hou et al. (2024). Second, to the best of our knowledge, our study represents the first application of multimodal GPTs (diagram + prompt) to the software development sub-task of SE, which in GenSE is generally carried out via natural language-based prompts. Third, we have discussed and illustrated with an empirical example that diagram-based prompting may be advantageous, in terms of information transfer efficiency, compared to pure natural language prompting; a result that aligns with the discussion in Tufte (2003). Fourth, by reverse engineering code into diagrams, we have contributed to the code summarisation sub-task of SE, which is again predominantly carried out via natural language summaries. Fifth, while there exist studies that investigated the production of PlantUML code from hand-drawn diagrams Conrardy and Cabot (2024); Wang et al. (2024), we believe this is the first study in which we cover multiple elements of the SE stack, from conceptual modelling — which includes static and dynamic analysis of the software system — to implementation. Finally, in addition, the use of existing (or new) design patterns in the context of Few-Shot prompting for software system design does not seem to have been investigated before in the literature, and hence merits attention in future studies.
Next, we shall look at which long standing issues in SE may be addressed by the use cases discussed in our work. In Baltes and Diehl (2014), the authors speculated that, as documentation is frequently poorly written and out of date, sketches may serve as a supplement to conventional documentation like source code comments. Consequently, they advocated the development of suitable tools to support archiving and retrieval of such sketches. No such tool has been developed ever since and poorly documented code remains a persistent issue in SE. However, GenAI has already started to change this picture via automatic generation of code comments. And yet, if sketches and diagram (both formal and informal) can be transformed, as we have shown, into a preliminary draft of the desired source code, then producing reliable visual descriptions of the software system under development suddenly becomes appealing. Likewise, if it is possible to selectively and cheaply reverse engineer part of such software system to visually inspect the behaviour of classes and methods, the developer becomes equipped with a formidable arsenal of high resolution communication tools to enhance code understanding, boost team communication, and resolve the long lasting conflict between investing time in developing working software or writing comprehensive documentation. These are all important and novel directions of enquiry that should be investigated in future studies.

### v3

@@ 2.1 Software Engineering and UML
The Unified Modelling Language (UML) is a general-purpose standardised visual language for designing systems. In SE, UML is utilised in three main areas: use case development, static analysis, and dynamic analysis of the software system. Use case development is a key step of requirement analysis. Use cases [^1] User stories in Agile Software Development Beck et al. (2001) describe how a user interacts with a system or product to achieve a specific goal. Static analysis of a software system describes the structure of a system by showing its classes, their attributes, operations (or methods), and the relationships among objects. Dynamic analysis expresses and model the behaviour of the system over time. More specifically, in UML, use case diagrams support use case development. Class diagrams support static analysis; and interaction (sequence, activity, collaboration) diagrams support dynamic analysis. However, it should be noted that UML is a vast language that finds countless applications to go beyond the three areas here considered.
@@ 2.2 Generative Artificial Intelligence
Prompts are instructions or input given to a large language model (LLM) to generate a specific response. Prompt engineering — the design prompting strategies to query language models — offers a cost-effective way to adapt pre-trained models without full fine-tuning.
Single Prompt Techniques. Zero-Shot Prompting involves providing tasks with natural language without additional context, relying on the model’s pre-existing capabilities. Few-Shot Prompting includes providing examples within the prompt to guide the model, enhancing its performance on complex tasks by exposing it to the input and output patterns Brown et al. (2020). Chain of Thought Prompting Wei et al. (2024) aids in breaking down reasoning tasks into smaller steps to improve outcome accuracy, using either zero-shot or few-shot methods to encourage step-by-step thinking.
Multiple Prompt Techniques. Voting/self-consistency Wang et al. (2023) involves generating multiple responses and selecting the most common result, which can improve accuracy, especially for complex reasoning tasks. Divide and Conquer methods split tasks into subtasks handled in sequence for improved manageability and precision, seen in Directional Stimulus Prompting Li et al. (2024a), Generated Knowledge Liu et al. (2021), and Prompt Chaining. Self-evaluation asks the model to verify output accuracy, exemplified by Reflexion Shinn et al. (2024) and Tree of Thoughts Yao et al. (2024), enabling iterative improvement.
Retrieval-Augmented Generation (RAG) Fan et al. (2024) and ReAct Yao et al. (2023) combine LLMs with external systems to improve context handling and output relevance.
@@ 2.3 The role of GenAI in Software Engineering
Code generation has been object of investigation for a long time in the AI community. Early works used symbolic and neural-semiotic approaches Alur et al. (2013). However, recent neurolinguistic models, such as GPT-4 Liu et al. (2024) and Copilot Ma et al. (2023), can generate code directly from natural language descriptions. While there are several works and benchmarks in the literature concerned with method-level code generation, to the best of our knowledge there is only one study and benchmark on class-level code generation Du et al. (2023), and none on diagram-level code generation. Moreover, none of the studies listed in the above surveys focus on code generation leveraging multimodal prompts that include sketches & diagrams.
@@ 3 Methodology
We develop a portfolio of novel GenSE use cases that, to the best of our knowledge, have not been previously investigated in the literature.
We utilise Microsoft Copilot in its web-based version, which is based on GPT-4o and allows image attachments. To ensure reproducibility of the discussion in Sections 4.1-4.4, we have also developed an equivalent pipeline in Gemini, formalised in a Jupyter Notebook based on gemini-1.5-flash, which is included in the supplementary material (SM). In both cases, we left all LLM parameters to their default settings and we did not specify any role or context instructions.
Table 1 maps use cases discussed in the rest of this work to relevant SE sub-tasks of interest. All use cases are based on a duly documented interaction with the LLMs that takes the form of a chat comprising multiple rounds of questions and responses (Prompt Chaining), which allow users to step incrementally towards answers and thus get help with multipart problems Google AI (2025). The core principle underpinning all our use cases is to illustrate possible strategies to leverage UML diagrams in order to guide the software development process. More specifically, in Section 4.1 we leverage class diagrams to guide the LLM in the context of implementing relevant classes, attributes, and operations; in Section 4.2 we leverage interaction diagrams to guide the LLM in the context of implementing the desired system behaviour; in Section 4.3 we leverage hand-drawn activity diagrams to guide the LLM in the context of implementing a given method; in Section 4.4 we leverage design patterns to influence the design of a given system. Finally, in Section 4.5 we present three additional case studies: in the first, we leverage hand-drawn diagrams and design patterns to implement a mathematical expressions evaluator; in the second and third, which feature a higher degree of complexity, we leverage class diagrams to implement a tic tac toe game and a game of checkers, respectively.
@@
dynamic modelling 4.2 SD/DE/CS
hand drawn diagrams 4.3 DE/CS
design patterns 4.4 SD
expression evaluator 4.5 SD/DE
tic tac toe 4.5 SD/DE
checkers 4.5 SD/DE
@@ 4 Multimodal GPTs in software development
The role of sketches and diagrams in the daily work of software developers has been investigated in Baltes and Diehl (2014). This study found that most practitioners produced informal hand-drawn diagrams and did not use UML; those using UML, tended to use it informally and selectively. We argue that this stems from the fact that the adoption of a “formal” notation, at present, bears no advantage with respect to an “informal” one. Over several decades, firms have repeatedly tried to develop tools (e.g. IBM Rational) that could automatically generate working code from UML diagrams, or reverse engineer UML diagram from existing code. While these technologies still exist, it is rare to find developers who routinely develop a complete set of UML diagrams and then translate this to code using such tools. The lacklustre success of these tools is likely due to a fundamental misunderstanding of what UML is: a language for capturing and communicating conceptual requirements, not one aimed at describing a complete system. In other words, UML diagrams are most useful when they are high level and sufficiently abstract. Generation of complete working code requires such diagrams to reach a level of detail that is equivalent to the working code itself; but generating diagrams at this level of abstraction would be a complete waste of time: why then not generating the code itself directly?
@@ 4.2 Dynamic modelling
We may now want to proceed and describe the dynamic behaviour of the system. Rather than using natural language, we may describe specific aspect of such behaviour by leveraging suitable UML interaction diagrams, such as the sequence diagram in Figure 3.
@@ 4.3 Hand-drawn diagrams
In this specific instance, Copilot returned a diagram with a small number of syntax errors. The errors were minor issues with the PlantUML syntax of the two while loops in the diagram, and could be easily fixed by hand. The resulting activity diagram is shown in Figure 9. Conversely, a second prompt asking Copilot to generate a diagram in PlantUML notation by translating directly the diagram in Figure 8 produced no errors. Since the focus of our discussion is facilitating — and not fully automating — SE activities, the presence of minor errors is of no concern, as these can easily be addressed in follow up prompts.
@@ 4.4 Design patterns
4.5 Implementation of complex systems
In this section, we discuss the implementation of three more complex applications. In all cases, the interaction takes the form of a chat comprising multiple rounds of questions and responses (Prompt Chaining), which allows the user to step incrementally towards a working implementation.
4.5.1 Evaluating mathematical expressions
In this first application, we consider a class diagram (provided in our SM) that conceptually captures a system that evaluates mathematical expressions comprising additions and subtractions. The system comprises five interrelated classes. We were able to obtain a complete implementation with a single multimodal prompt (diagram + instruction to implement the diagram). An additional prompt was required to implement an evaluator by leveraging the Visitor design pattern.
4.5.2 Tic tac toe
We consider the class diagram [^3] https://github.com/pelensky/JavaTTT/blob/master/UML.pdf for a tic tac toe game. The diagram comprises over twenty interrelated classes. We use this diagram to guide the development of a tic tac toe Python application. The detailed interaction with Copilot is illustrated in our SM. This interaction involves seven prompts in which Copilot is asked to fully implement specific classes in the diagram. In line with traditional SE practices, we start from classes featuring low complexity and connectivity, and we then move towards other classes that depend on those already implemented. This is followed by two prompts to correct errors encountered while trying to execute the application. In each of these prompts the full stack trace of the error is fed to Copilot. Finally, an additional prompt is used to tailor board visualisation. A complete application that neatly matches the conceptual model presented in the class diagram is hence obtained in ten short prompts comprising a single sentence each.
4.5.3 Checkers
We consider the class diagram [^4] https://app.genmymodel.com/api/repository/wanex505/checkers for a checkers game. The diagram comprises over ten interrelated classes and a complex underpinning logic. We use this diagram to guide the development of a checkers Python application. The detailed interaction with Copilot is illustrated in our SM. This interaction involves nine prompts in which Copilot is asked to fully implement specific classes in the diagram. This is followed by an additional prompt to develop a text-based drawing logic. Ten additional prompts are then required to correct errors encountered while trying to execute the application and play the game. These errors include both exceptions (e.g. a class attribute is missing) as well as conceptual errors that affect the game logic (a piece is not removed from the board after being jumped over) or the user interaction (e.g. the user is requested to input a move, but it is not stated if they are playing as black or white). A complete application that matches the conceptual model presented in the class diagram is hence obtained in twenty short prompts comprising a single sentence each.
In our SM we provide a more detailed description of each case study, as well as Jupyter Notebooks reporting the complete interactions with Copilot.
@@ 5 Discussion and conclusions
We shall next focus on the research gaps we bridged in GenSE. First, our study addresses the lack of applications of GenAI to the software design sub-task of SE, a literature gap identified in Huang et al. (2024); Hou et al. (2024). Second, to the best of our knowledge, our study represents the first application of multimodal GPTs (diagram + prompt) to the software development sub-task of SE, which in GenSE is generally carried out via natural language-based prompts. We argue that diagram-based prompting may be advantageous, in terms of information transfer efficiency, compared to pure natural language prompting, as diagrams leverage background knowledge associated with their visual elements and structure, thereby leading to compression Li et al. (2024b). Third, by reverse engineering code into diagrams, we have contributed to the code summarisation sub-task of SE, which is again predominantly carried out via natural language summaries. Fourth, while there exist studies that investigated the production of PlantUML code from hand-drawn diagrams Conrardy and Cabot (2024); Wang et al. (2024), we believe this is the first study in which we cover multiple elements of the SE stack, from conceptual modelling — which includes static and dynamic analysis of the software system — to implementation. Moreover, to the best of our knowledge, this is the first study that focus on diagram-level (rather than method-level or class-level) code generation. Finally, the use of existing (or new) design patterns in the context of Few-Shot prompting for software system design does not seem to have been investigated before in the literature, and hence merits attention in future studies.
Next, we shall look at which long standing issues in SE may be addressed by the use cases discussed in our work. In Baltes and Diehl (2014), the authors suggested that sketches could supplement often outdated and poorly written documentation, advocating for tools to archive and retrieve these sketches. No such tool exists and poorly documented code remains a persistent issue in SE. However, GenAI has already started to change this picture via automatic generation of code comments. And yet, if sketches and diagram (both formal and informal) can be transformed, as we have shown, into a preliminary draft of the desired source code, then producing reliable visual descriptions of the software system under development suddenly becomes appealing. Likewise, if it is possible to selectively and cheaply reverse engineer part of such software system to visually inspect the behaviour of classes and methods, the developer becomes equipped with a formidable arsenal of high resolution communication tools to enhance code understanding, boost team communication, and resolve the long lasting conflict between investing time in developing working software or writing comprehensive documentation. These are all important and novel directions of enquiry that should be investigated in future studies.